OpenStudy (anonymous):

(x^2y)/(3y^3x^3) times (18x^4y^2)/(xy^6)

OpenStudy (anonymous):

help!

OpenStudy (anonymous):

\[\frac{x^2y}{3y^3x^3}*\frac{18x^4y^2}{xy^6}\] \[\frac{6x^2}{y^6}\]

OpenStudy (anonymous):

help!

OpenStudy (anonymous):

can you show me your work? i do not know how you got to that answer

OpenStudy (anonymous):

\[\frac{x^2\cancel{y}}{\cancel{3y^3x^3}}*\frac{6*\cancel{3x^4y^2}}{\cancel{x}y^6}=\frac{6x^2}{y^6}\]